### v4

# Extending edge colorings of distance-3 matchings in the Cartesian product of graphs

Abstract
We investigate the problem of extending partial edge colorings in Cartesian products of graphs, with a particular focus on cases where the precolored edges form a matching. Casselgren, Granholm, and Petros conjectured that any precolored distance-3 matching in $G=C^{d}_{2k}$ can be extended to a $2d$-edge coloring. In this paper, we prove a theorem that implies this conjecture. Especially, our main result establishes that a precolored distance-3 matching in the Cartesian product of certain class 1 graphs can be extended to an edge coloring that uses at most as many colors as the chromatic index, provided that certain degree conditions are satisfied. In the second part of the paper, we extend these results to Cartesian products of other types of graphs as well.
Keywords: Precoloring extension; Edge coloring; Bipartite graph; Cartesian product
1 Introduction
In this paper, we deal with proper edge colorings of graphs. Throughout the paper, instead of proper edge coloring, we often say just coloring. The edge chromatic number (or chromatic index) of a graph $G$ is denoted by $\chi^{\prime}(G)$. According to Vizing’s Theorem [12], if $G$ is a simple graph with maximum degree $\Delta(G)$, then its chromatic index is either $\Delta(G)$ or $\Delta(G)+1$. Graphs with $\chi^{\prime}(G)=\Delta(G)$ are called Class 1 graphs. (In particular, bipartite graphs are proved to be Class 1 graphs.) An (edge) precoloring (or partial edge coloring) of a graph $G$ is a proper edge coloring of some edge set $E_{0}\subseteq E(G)$.
Definition 1.1.
We call a precoloring of the edges in $E_{0}$ extendable if there is a proper edge coloring with $\chi^{\prime}(G)$ colors such that the color of the edges in $E_{0}$ are the prescribed colors. Such a coloring is called the extension of the precoloring.
We recall that the problem of extending a given edge precoloring is an NP-hard problem, even for 3-regular bipartite graphs. [8], [10]
Definition 1.2.
The Cartesian product $G_{1}\square G_{2}$ of graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ is a graph such that the vertex set of $G_{1}\square G_{2}$ is the Cartesian product $V(G_{1})\times V(G_{2})$ and two vertices $(u_{1},u_{2})$ and $(v_{1},v_{2})$ are adjacent in $G_{1}\square G_{2}$ if and only if either
  • $u_{1}=v_{1}$ and $u_{2}$ is adjacent to $v_{2}$ in $G_{2}$ or
  • $u_{2}=v_{2}$ and $u_{1}$ is adjacent to $v_{1}$ in $G_{1}$.
Notation 1.3.
$G^{d}$ denotes the $d$th power of the Cartesian product of $G$ with itself.
Notation 1.4.
For the edges $xy=e\in E(G)$ and $pq=f\in E(H)$, $xy\square pq=e\square f$ denotes the subgraph (a 4-cycle) induced in $G\square H$ by $\{(x,p),(x,q),(y,p),(y,q)\}$.
One of the basic concepts in the paper is as follows.
Definition 1.5.
(The brick-neighborhoods of an edge in the Cartesian product) Let $yz\in E(G)$ be an edge. Then take a path $xyzw$ in $G$ and an edge $pq\in E(H)$. Then a brick-neighborhood of the edge $(y,q)(z,q)$ is the Cartesian product of the path $xyzw$ and the edge $pq$ ($xyzw\square pq$). Symmetrically, we can define brick-neighborhoods starting with an edge in $H$ (i.e., we take a 3-path in $H$ and an edge in $G$).
The edge $(y,p)(z,p)$ and the edges in the brick-neighborhood incident to $(y,q)$ or $(z,q)$ (altogether six edges) are called internal edges, the other (four) edges of the brick-neighborhood are called external edges.
Remark 1.6.
Notice that an edge $e\in E(G\square H)$ typically has several brick-neighborhoods depending on the choice of the edges $xy,zw$ and $pq$.
Definition 1.7.
Let $G$ be a graph, and let us prescribe some color for some edges of $G$. We call a (proper edge) coloring of $G$ with $\chi^{\prime}(G)$ colors admissible if it is an extension of the precoloring.
Problem 1.8.
Let $G$ and $H$ be bipartite (or Class 1) graphs with maximum degrees $\Delta(G)$ and $\Delta(H)$, respectively. We would like to color the edges of the Cartesian product $G\square H$ with $\Delta(G)+\Delta(H)$ colors, even if we prescribe the color of some edges. Under what conditions does the graph have an admissible coloring?
A natural condition is that the precolored edges are far from each other. The problem of extending a partial edge coloring with distance constraints on the precolored edges goes back to the work of Albertson and Moore [1].
Definition 1.9.
In a connected graph, the distance $d(x,y)$ between vertices $x$ and $y$ is the length of the shortest $xy$ path and the distance $d(e,f)$ of edges $e=xy$ and $f=zw$ is $\min\{d(x,z),d(x,w),d(y,z),d(y,w)\}$.
Definition 1.10.
The matching $M$ is a distance-k matching, if for every $m_{1},m_{2}\in M:d(m_{1},m_{2})\geq k$.
Casselgren, Granholm, and Petros [2] proved
Theorem 1.11.
If $E_{0}\subseteq E(G)$ is a distance-4 matching of $G=C^{d}_{2k}$, then any $2d$-coloring of $E_{0}$ can be extended to a proper $2d$-edge coloring of $G$.
Some related theorems are proved in [3], [4], [5], [6], [9] and [11].
In the paper [2], the authors raised the conjecture as follows.
Conjecture 1.12.
If $E_{0}\subseteq E(G)$ is a distance-3 matching of $G=C^{d}_{2k}$, then any $2d$-coloring of $E_{0}$ can be extended to a proper $2d$-edge coloring of $G$.
Remark 1.13.
It is easy to give an example of a graph in which the minimal distance of the precolored edges is at least two, but does not have an admissible coloring (e.g., see [2]).
Remark 1.14.
It is clear that Theorem 1.11 also holds for the Cartesian product $G\square H$ of arbitrary Class 1 graphs $G$ and $H$, instead of $C^{d}_{2k}$. We may assume that $G$ and $H$ are regular graphs (see Lemma 2.2), and under this assumption, the proof of Theorem 1.11 applies directly.
We recall that the direct product of a Class 1 graph with any other graph is also a Class 1 graph. Therefore, in the case of a product with multiple factors, if at least one of the factors is Class 1, then the entire product is Class 1 as well.
The main theorem of this paper is the following theorem that implies Conjecture 1.12.
Theorem 1.15.
Let $k$ be an integer with $k\geq 2$, and let $G_{i}$ be Class 1 graphs for each $i\in\{1,\dots,k\}$, with maximum degree $\Delta(G_{i})$, respectively. Suppose that some edges of the Cartesian product $G_{1}\square G_{2}\square\cdots\square G_{k}$ are precolored using at most $\sum_{i=1}^{k}\Delta(G_{i})$ colors, ensuring that the distance between any two precolored edges is at least three. If $\sum_{i=1}^{k}\Delta(G_{i})$ is even and satisfies $2\Delta(G_{i})<\sum_{i=1}^{k}\Delta(G_{i})\forall i\in\{1,\dots,k\}$, then this precoloring is extendable.
Remark 1.16.
Notice that if the distance between any two precolored edges is at least three, then the edge set of two brick-neighborhoods of different precolored edges are either edge-disjoint or share an external edge.
Remark 1.17.
Theorem 1.15 solves Problem 1.8 in the case where $\Delta(G)=\Delta(H)$ and also Conjecture 1.12, as the degree conditions are clearly satisfied in both cases.
In the second section, we also study the problem in the case of the Cartesian product of a bipartite graph and an edge. Galvin’s theorem [7] is used in the proof of this statement, so the proof does not work for the Cartesian product of a Class 1 graph and an edge. However, if the List Coloring Conjecture holds for Class 1 graphs, then the proof automatically extends to this case as well. The final section presents some related results on Cartesian products with odd cycles. The following definition will be useful in the proofs.
Definition 1.18.
(Canonical coloring of the Cartesian product of graphs) Let us fix a proper coloring of each Class 1 graph $G_{i}$ for $i\in\{1,\dots,k\}$ using $\Delta(G_{i})$ colors, respectively, resulting in a total of $\sum_{i=1}^{k}\Delta(G_{i})$ distinct colors. We define the canonical coloring of the graph $G=G_{1}\square G_{2}\square\cdots\square G_{k}$ as follows. If an edge $(x_{1},\dots,x_{k})(y_{1},\dots,y_{k})\in E(G)$ connects two vertices that differ in the $j$th coordinate, we assign it the same color as the edge $x_{j}y_{j}$ in $G_{j}$.
The canonical coloring is a proper coloring with $\sum_{i=1}^{k}\Delta(G_{i})$ colors, and we apply certain local operations to make it admissible.
2 Admissible edge-colorings of the Cartesian
product of bipartite graphs
In this section, we prove Theorem 1.15. The basic idea of the proof is an improved and adapted version of the proof of Theorem 5.1 in [2] on extending partial edge colorings of iterated Cartesian products of cycles and paths.
Lemma 2.1.
Given $k\geq 2$ disjoint sets $H_{1},\dots,H_{k}$ and their union $H=H_{1}\cup\dots\cup H_{k}$ where $|H|$ is even and satisfies $2|H_{i}|<|H|$ for all $i\in\{1,\dots,k\}$. We can pair off the elements of $H$ so that no two elements in any pair belong to the same set $H_{i}$.
Proof.
The proof proceeds by induction on $|H|$. For $|H|<8$, the statement holds trivially. Now, assume that the statement holds for all sets $H^{\prime}$ of size smaller than $8$, and consider a set $H$ that satisfies the given conditions. Select one element from each of the two largest sets and pair them together. After this pairing, the remaining sets still satisfy the conditions: The size of $|H|$ decreases by $2$, so it remains even. The sizes of the two largest sets decrease by $1$ each, which means that their doubled sizes decrease by $2$. The third largest set could not have had more than $\frac{|H|-2}{2}$ elements before pairing, because otherwise the total size of $H$ would be at least $3\frac{|H|-2}{2}$, which exceeds $|H|$, contradicting the assumption that $|H|\geq 8$. Thus, after removing the paired elements, the remaining elements still satisfy the conditions of the statement and we can complete the pairing. ∎
Lemma 2.2.
Let $k$ be an integer with $k\geq 2$, and let each $G_{i}$ (for $i\in\{1,\dots,k\}$) be a Class 1 graph. By adding extra vertices and edges to each graph $G_{i}$, we can make them $\Delta(G_{i})$-regular while preserving their Class 1 property and ensuring that if the distance between any two edges in the Cartesian product of the original graphs $G_{i}$ is at least 3, then the same holds in the Cartesian product of the regularized version as well.
Proof.
In the proof, we will show that a single factor can be made regular. This procedure can then be applied individually to each factor, thereby proving the statement.
We describe a possible procedure for adding these extra vertices and edges to a graph $H$. Starting with $H$, whenever a vertex has a degree less than $\Delta(H)$, we duplicate the entire graph and connect each such vertex to its counterpart in the copy. Duplicating $H$ ensures that any vertex of degree less than $\Delta(H)$ gains exactly one new neighbor, increasing its degree without exceeding $\Delta(H)$. This operation increases the minimum degree without raising the chromatic index. We repeat this process until the graph becomes regular.
It is clear that the graph remains Class 1 at every step, since we can color both copies identically. Furthermore, since new edges are only added between vertices whose degrees are still below $\Delta(H)$, we can assign proper colors to these new edges without increasing the total number of colors used.
Let $G$ and $H$ be Class 1 graphs, and let $H^{*}$ denote the regular extension of $H$ obtained by this process. If two vertices had distance at least 3 in the graph $G\square H$, then their distance will also be at least 3 in the graph $G\square H^{*}$.
Distance one is clearly impossible, since no new edges are added within the $G\square H$ subgraph. Distance two is also not possible, as this would require one of the following two configurations:
  • In the first case, the three vertices are $(u,v)$, $(u,w)$, and $(u,z)$, where $w$ must necessarily be a new vertex. Otherwise, this would already be a path of length two in $G\square H$, contradicting our assumption. However, based on the construction, each new vertex has at most one neighbor in the original graph, making this configuration impossible.
  • In the second case, the three vertices are $(u,v)$, $(u,w)$, and $(y,w)$. But in this case neither $u$ nor $w$ can be a new vertices, again contradicting the rules of the construction. Hence, this configuration is also impossible.
Proof of Theorem 1.15.
We can color the components of the graph one by one, so we may assume that the graph is connected. Let $G=G_{1}\square\dots\square G_{k}$ be used in the following. Consider a canonical coloring of the graph $G$. This is a proper coloring with $\Delta(G)$ colors, but is usually not an admissible coloring. Now, we modify the coloring step by step, preserving the number of colors and ensuring that the coloring is proper. If a precolored edge has the correct color, then we will not change it. If the prescribed color of a precolored edge is not its color in the canonical coloring, then we will correct it by some local operations. So, we finally get an admissible coloring.
We may assume that the graphs $G_{i}$ are regular according to Lemma 2.2. Notice that if this new graph has an admissible coloring, then its restriction for $G$ is an admissible coloring, too.
For the edges $e\in E(G_{i})$ and $f\in E(G_{j})$, the subgraph $e\square f$ of $G$ is called a square. In the canonical coloring, the edges of each square are colored with two colors. A rotation in a 2-colored square is the swap of the colors of its edges. (On the other edges, the coloring is not changed.) After the rotation in a square, the number of colors does not change and the coloring remains proper. Since in each step we perform a sequence of rotations, these two constraints (the number of colors is $\Delta(G)$ and the coloring is proper) will be fulfilled throughout. So we pay attention to three things:
  1. 1.
    Only rotate a square with opposite edges of the same color. (The rotation is defined on 2-colored squares only!)
  2. 2.
    Do not change the color of a precolored edge if it has its prescribed color.
  3. 3.
    The color of every precolored edge should be the prescribed color after some steps.
STEP 1: Suppose that an edge $(x_{1},\dots,x_{k})(y_{1},\dots,y_{k})\in E(G)$ connects two vertices that differ in the $j$th coordinate, and its prescribed color appears in the coloring of the graph $G_{l}$ for some $l\neq j$. Since $G_{l}$ is regular, there must be an edge $(x_{l},z)\in E(G_{l})$ that got this prescribed color in the canonical coloring. In this case, we perform a rotation on the square $x_{j}y_{j}\square x_{l}z$. This rotation affects only the internal edges of a brick-neighborhood of the precolored edge, ensuring that the brick-neighborhoods of the other precolored edges remain unchanged (see Remark 1.16).
STEP 2: Suppose that an edge $(x_{1},\dots,x_{k})(y_{1},\dots,y_{k})\in E(G)$ connects two vertices that differ in the $j$th coordinate, and its prescribed color appears in the coloring of the graph $G_{j}$, but the color of this edge in the canonical coloring is different from the prescribed color. Then there are edges $sx_{j},y_{j}t\in E(G_{j})$, which have this color in the canonical coloring. Now, take an arbitrary edge $(x_{i},w)$ of the graph $G_{i}$ for some $i\neq j$ and perform a rotation sequentially on the squares $sx_{j}\square x_{i}w$, $y_{j}t\square x_{i}w$ and finally on $x_{j}y_{j}\square x_{i}w$.
In this case, the coloring changes on the external edges, as well. Therefore, we aim to select one of the brick-neighborhoods of each of these type precolored edges in such a way that they are pairwise edge-disjoint. If we can do this, then we can achieve an admissible coloring by applying the second type of step above to the brick-neighborhoods chosen in this way.
The degrees of the graphs $G_{1},\dots,G_{k}$ (and consequently the colors used) satisfy the conditions of Lemma 2.1 Therefore, the colors used for the coloring $G_{1},\dots,G_{k}$ can be paired so that no two colors in any pair originate from the same graph $G_{i}$. For a precolored edge $yz$, we select the brick-neighborhood $B=xyzw\square pq$ such that the color of $xy$ and $zw$ is the prescribed color of $yz$ and the color of $pq$ is the pair of this color. Let $f(B)$ denote this pair of colors, i.e., the colors assigned to the edges of the end squares $xy\square pq$ or $zw\square pq$. For any two chosen brick-neighborhoods $B_{1}$ and $B_{2}$ (corresponding to precolored edges), sets $f(B_{1})$ and $f(B_{2})$ are either identical or disjoint.
Notice that any two selected brick-neighborhoods are edge-disjoint. Suppose, for contradiction, that two brick-neighborhoods share an external edge (by Remark 1.16) In that case, they must have two end squares sharing exactly one edge, leading to two color pairs that share exactly one color, which contradicts our pairing condition. This completes the proof of Theorem 1.15. ∎
Remark 2.3.
The above proof for two graphs relies on the equality $\Delta(G)=\Delta(H)$. We believe that the statement holds in general as well. It would be a promising step to prove the theorem at least for some values of $\Delta(G),\Delta(H)$ that are not equal. A natural first step is to consider cases where $\Delta(G)$ and/or $\Delta(H)$ is small. The following theorem settles the very first case $\Delta(H)=1$.
Theorem 2.4.
Let $G$ be a bipartite graph with maximum degree $\Delta(G)$. Suppose that the color of some edges of the Cartesian product $G\square K_{2}$ is prescribed using at most $\Delta(G)+1$ colors such that the distance between any two precolored edges is at least three. Then, this precoloring is extendable.
Proof.
Let $V(K_{2})=\{a,b\}$. To prove the theorem, we will consider a list edge-coloring problem for (a subgraph of) $G$. First, assign the list of colors $\{1,2,...,\Delta(G)+1\}$ to every edge of $G$. If an edge $(u,a)(v,a)$ or $(u,b)(v,b)$ has a prescribed color (note that both cannot be precolored simultaneously), then we remove the edge $uv$ from $G$ and remove the prescribed color from the lists of all edges adjacent to $uv$. If there is a prescription on an edge $(u,a)(u,b)$, then we remove the prescribed color from the list of all edges incident to $u$.
The resulting graph $G^{\prime}$ remains bipartite, with a maximum degree of at most $\Delta(G)$. Moreover, each color list still contains at least $\Delta(G)$ colors since the distance condition ensures that at most one color was removed from each list. By Galvin’s theorem [7], $G^{\prime}$ admits a proper list-coloring from these lists. This coloring of $E(G^{\prime})$ can then be extended by incorporating the prescribed colors to obtain a coloring of $E(G)$ in both copies of $G$ within the Cartesian product.
Since the set of colors assigned to the edges incident to $u$ is the same in both copies, there is always a remaining color available to color the edge $(u,a)(u,b)$. If $(u,a)(u,b)$ had a prescribed color, the remaining available color must be the prescribed one, ensuring that the final coloring is admissible. ∎
Remark 2.5.
A similar proof also works for the Cartesian product $C_{2k+1}\square K_{2}$, we can greedily color the appropriately precolored $C_{2k+1}$ with $3$ colors instead of using Galvin’s theorem.
By repeated applications of Theorem 2.4 we get the corollary as follows.
Corollary 2.6.
Let $G$ be a bipartite graph with maximum degree $\Delta(G)$. Suppose that the color of some edges of the Cartesian product $G\square K_{2}^{\alpha}$ (especially for $G\square C_{4}$) is prescribed using at most $\Delta(G)+\alpha$ colors, so that the distance between any two precolored edges is at least three. Then this precoloring is extendable.
Notice that Corollary 2.6 implies Conjecture 1.12 for $k=2$.
3 Admissible edge-colorings of the Cartesian
product $C_{2k+1}\square C_{2l+1}$
Conjecture 1.12 is about the product of even cycles. A natural continuation of this is to also examine the extendability of precolorings in the case of products of odd or odd and even cycles. We can settle the case of two odd cycles.
Theorem 3.1.
Suppose that the color of some edges of the Cartesian product $C_{2k+1}\square C_{2l+1}$ is prescribed using at most $5$ colors such that the distance between any two precolored edges is at least three. Then this precoloring is extendable.
Remark 3.2.
This is the strongest possible statement. The graph $C_{2k+1}\square C_{2l+1}$ has an odd number of vertices, which means that it does not contain perfect matching. Therefore, it cannot be 4-edge-colorable, as that would require its edges to be divided into four perfect matchings. It is easy to give a non-extendable precoloring of the graph $C_{2k+1}\square C_{2l+1}$ where the minimum distance between precolored edges is at least two (see Remark 1.13 or [2] ).
Proof of Theorem 3.1.
Notation 3.3.
[…]
It is easy to see that there is a square that is not part of the brick-neighborhood of any precolored edge. W.l.o.g., we may assume that $(2k,2l)(2k,0)$ $(0,0)(0,2l)$ is such a square.
Choose an arbitrary color from the five available ones. Color the edges that are not of the type $(i,2l)(i,0)$ or $(2k,j)(0,j)$ using the remaining four colors, taking into account the prescriptions (unless the fifth color is prescribed). We can do this according to Theorem 1.15. This subgraph that we color in this first step will be referred to as a truncated grid from now on. It is important to note that we color this truncated grid using the method used in the proof of Theorem 1.15, which we will use later in the proof.
Afterward, the edges of types $(i,2l)(i,0)$ and $(2k,j)(0,j)$ are colored with the fifth color. (This will not be a proper coloring, but aside from the edges of the square $(2k,2l)(2k,0)(0,0)(0,2l)$, it is proper.) If an edge $(i_{1},j_{1})(i_{2},j_{2})$, where $\{i_{1},j_{1},i_{2},j_{2}\}\cap\{0,2k,2l\}=\emptyset$, has the fifth color prescribed, it can be colored with this color and no further changes are needed. If an edge $(i_{1},j_{1})(i_{2},j_{2})$ has exactly one of $i_{1},j_{1},i_{2},j_{2}$ from the set $\{0,2k,2l\}$ and has the fifth color prescribed, it is adjacent to exactly one edge of this color. In this chase, we can swap the color of these two edges. Similarly, if an edge of type $(i,2l)(i,0)$ or $(2k,j)(0,j)$ has a prescription such that it has exactly one adjacent edge of this color, then we can swap the color of these two edges. If an edge of type $(i,2l)(i,0)$ or $(2k,j)(0,j)$ has a prescription such that it has two adjacent edges of this color, then they define a square, and we can rotate on this square. If there is a prescription of the fifth color on an edge of type $(i,*)(j,*)$ or $(*,i)(*,j)$ (the symbol $*\in\{0,2k,2l\}$ denotes the same value in the two vertices), then there is a square with two sides colored with the fifth color; we can rotate on this square.
All these modifications result in an admissible coloring, except for the edges of the square $(2k,2l)(2k,0)(0,0)(0,2l)$. After the coloring of the truncated grid, we only change the coloring on the internal edges of the brick-neighborhoods.
Finally, we need to fix the coloring on the square $(2k,2l)(2k,0)(0,0)(0,2l)$. In the canonical coloring of the truncated grid, the two paths should be colored using the colors $1,2$ and $3,4$. The edges $(2k,2l)(2k,0)$ and $(0,2l)(0,0)$ are colored with the fifth color. At least one pair of opposite edges of the square $(2k,2l)(2k,0)(0,0)(0,2l)$ remains colored with $5$, w.l.o.g., we may assume that these two edges are like this.
Now, consider the squares $A=(2k-1,0)(2k,0)(2k,1)(2k-1,1)$ and $B=(0,0)(0,1)(1,1)(1,0)$. If there exists a color other than $5$ that is not used in the coloring of either $A$ or $B$, then we can assign that color to the edge $(2k,0)(0,0)$. In fact, we only need to check the colors of the edges incident to the vertices $(2k,0)$ or $(0,0)$.
In the canonical coloring, the colors $2$ and $3$ appear in the coloring of the square $A$, while the colors $1$ and $3$ appear in the coloring of the square $B$. If these colors remain unchanged, we can color the edge $(2k,0)(0,0)$ with the color $4$. However, this situation changes only if at least one of the edges $(2k-1,0)(2k-1,1)$, $(2k-1,1)(2k,1)$, $(2k,1)(0,1)$, $(0,1)(1,1)$ and $(1,1)(1,0)$ is precolored.
If the edges $(2k-1,0)(2k,0)$ and $(2k,0)(2k,1)$ or $(0,0)(0,1)$ and $(0,0)(1,0)$ are external edges in the brick-neighborhood of a precolored edge, this will not affect the overall coloring. During rotations, the colors of these edges may change, but they will still have the same two colors, just in a different order.
In most cases, the coloring of the squares $A$ and $B$ remain unchanged throughout the modifications, allowing the edge $(2k,0)(0,0)$ to be assigned the color $4$. From this point on, we will only consider the cases where deviations from this standard situation occur. Due to symmetry and the distance condition for precolored edges, it is sufficient to examine the following cases:
  • The edge $(2k,1)(0,1)$ is prescribed to be color $3$. In this case, we perform a rotation on the square between $A$ and $B$, namely $(2k,0)(2k,1)(0,1)(0,0)$. As a result, we change the color of the edge $(2k,0)(2k,1)$ to $1$ and the color of the edge $(0,0)(0,1)$ to $2$. Consequently, the color of the edge $(2k,0)(0,0)$ becomes $3$.
  • The edge $(2k-1,1)(2k,1)$ is prescribed to be color $1$, and we performed the rotations in the brick-neighborhood containing $A$. As a result, the square $A$ will be colored with the colors $1$ and $2$, while the square $B$ remains unchanged, colored with the colors $1$ and $3$. Therefore, the edge $(2k,0)(0,0)$ can be colored with $4$.
  • The edge $(2k-1,0)(2k-1,1)$ is prescribed to be color $4$, and we performed the rotations in the brick-neighborhood containing $A$, with no changes occurring in the square $B$. As a result, the square $A$ will be colored with colors $3$ and $4$, while the square $B$ remains unchanged, colored with the colors $1$ and $3$. Therefore, the edge $(2k,0)(0,0)$ can be colored with $2$.
  • The edge $(2k-1,0)(2k-1,1)$ is precribed to be color $4$, and we performed the rotations in the brick-neighborhood containing $A$. Similarly, the edge $(1,0)(1,1)$ is prescribed to be color $4$, and we performed the rotations in the brick-neighborhood containing $B$. As a result, both square $A$ and $B$ will be colored with the colors $3$ and $4$. Therefore, the edge $(2k,0)(0,0)$ can be colored with $2$.
Thus, we can fix the color of the edge $(2k,0)(0,0)$, and by applying the same method, we can also fix the color of the edge $(2k,2l)(0,2l)$. Consequently, we have completed the coloring in an admissible manner, thereby proving the statement. ∎
4 Concluding remarks
Several interesting problems remain open. The most natural question is whether the main theorem holds for two graphs even if $\Delta(G)\neq\Delta(H)$, or in other words, whether the theorem holds without the degree conditions.
Conjecture 4.1.
Let $G$ and $H$ be Class 1 (or bipartite) graphs with maximum degrees $\Delta(G)$ and $\Delta(H)$, respectively. Suppose that the color of some edges of the Cartesian product $G\square H$ is prescribed using up to $\Delta(G)+\Delta(H)$ colors such that the distance between any two precolored edges is at least three. Then this precoloring is extendable.
We studied the extendability of precolorings on the product of even cycles and on the product of odd cycles; however, the case of the product of an even cycle and an odd cycle remains unsolved.
Conjecture 4.2.
Suppose that the color of some edges of the Cartesian product $C_{2k}\square C_{2l+1}$ is prescribed using at most $4$ colors such that the distance between any two precolored edges is at least three. Then this precoloring is extendable.
Another interesting question is whether Theorem 2.4 holds for any graph $G$.
Conjecture 4.3.
Let $G$ be a graph with maximum degree $\Delta(G)$. Suppose that the color of some edges of the Cartesian product $G\square K_{2}$ is prescribed using at most $\Delta(G)+1$ colors such that the distance between any two precolored edges is at least three. Then, this precoloring is extendable.
Statements and declarations
The first author is partially supported by the Counting in Sparse Graphs Lendület Research Group. The research of the second author was partially supported by the National Research, Development and Innovation Office NKFIH. The authors declare that they have no conflict of interest. Data sharing is not applicable to this article, as no datasets were generated or analyzed during the current study.